OpenStudy (anonymous):

If the equation of a quadratic is y = 6x^2 -4x - 1, then the equation of the axis of symmetry is: A. B. C. 3 D. -3

OpenStudy (anonymous):

A. x= -1/3 B. x= 1/3

OpenStudy (shubhamsrg):

The parabola will be symmetrical for that co-ordinate of x where the derivative is 0. Have you learnt differentiation? maxima-minima to be more precise. ?